# Parity Quantum Computing as YZ-Plane Measurement-Based Quantum Computing

### v1

@@ Abstract
We show that universal parity quantum computing employing a recently introduced constant depth decoding procedure is equivalent to measurement-based quantum computation (MBQC) on a bipartite graph using only YZ-plane measurements. We further show that any unitary MBQC using only YZ-plane measurements must occur on a bipartite graph.
@@ I Background
In the measurement-based regime, computation is driven by single-qubit projective measurements restricted to the $XY$-, $XZ$- and $YZ$-planes of the Bloch sphere. A given computation is defined by one specific outcome for each measurement, and the restriction to the given planes allows for the correction of undesired outcomes via an effective application of an appropriate stabiliser element (or products thereof). However, even with these restrictions not every sequence of measurements for a given graph states is possible. The combination of graph state and measurements that do allow for computation are well characterised by a property called gflow (see Appendix A) which is known to be a necessary and sufficient condition for deterministic MBQC [22].
@@ II Results
In light of the above, it is prudent to demarcate the parity computing regime with respect to the MBQC regime. One could reasonably ask if there exist YZ-only measurement-based computations on graphs that are not bipartite. However, the following theorem demonstrates that this in fact not the case. The theorem also takes care of any issues regarding correction of measurements (this is explained further in Appendix A).
@@
The proof makes use of technical lemmas related to gflow which are proved along with the theorem in Appendix A.
@@ III Discussion
There are a number of consequences of these results worth mentioning already here. In [21], it was noted that the parity measurement-based encoding and decoding procedures can be implemented in constant depth regardless of architecture size. Since the decoding procedure corresponds to measuring vertices in one partition of a bipartite graph, it is clear that all measurements can be done simultaneously and corrected for in the other partition. The encoding procedure can be understood as measuring ancilla vertices of a larger graph state in the $X$-basis, which in particular produces the required bipartite graph (see e.g., [10, 11] for a characterisation of graph state deformations under Pauli measurements). It is known that all Pauli-measurements can be performed simultaneously in MBQC [18].
Bipartite graph states have previously received substantial attention in the MBQC literature; see e.g., [26, 27, 28]. The set of bipartite graph states includes both cluster states and GHZ states, which are both known to be useful for computation and communication protocols. There exist multi-particle entanglement purification protocols for bipartite graph states which exhibit favourable error thresholds for realistic scenarios [26, 27, 28]. Similar techniques may be of benefit for error mitigation in the parity regime - an avenue for future work.
Acknowledgements: We would like to thank B. Klaver and A. Messinger for useful discussions. This research was funded in whole or in part by the Austrian Science Fund (FWF) through the DK-ALM W1259125912591259-N27272727 and the SFB BeyondC F7102. For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. This work was also co-funded by the European Union (ERC, QuantAI, Project No. $101055129$). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.

### v2

@@ Abstract
We show that universal parity quantum computing employing a recently introduced constant depth decoding procedure is equivalent to measurement-based quantum computation (MBQC) on a bipartite graph using only YZ-plane measurements. We further show that any unitary MBQC using only YZ-plane measurements must occur on a bipartite graph. These results have a number of consequences and open new research avenues for both frameworks.
@@ I Background
In the measurement-based regime, computation is driven by single-qubit projective measurements restricted to the $XY$-, $XZ$- and $YZ$-planes of the Bloch sphere. A given computation is defined by one specific outcome for each measurement, and the restriction to the given planes allows for the correction of undesired outcomes via an effective application of an appropriate stabiliser element (or products thereof). However, even with these restrictions not every sequence of measurements for a given graph states is possible. The combination of graph state and measurements that do allow for computation are well characterised by a property called gflow which is known to be a necessary and sufficient condition for deterministic MBQC [22] (see [24] for the definition).
@@ II Results
In light of the above, it is prudent to demarcate the parity computing regime with respect to the MBQC regime. One could reasonably ask if there exist YZ-only measurement-based computations on graphs that are not bipartite. However, the following theorem demonstrates that this in fact not the case. The theorem also takes care of any issues regarding correction of measurements (see [24] for more details).
@@
The proof makes use of technical lemmas related to gflow which are proved along with the theorem in the Supplemental Material [24].
@@ III Discussion
As this work connects two previously distinct bodies of research, there are a number of consequences worth mentioning here. Firstly, our results provide insight into recent research in the parity framework. In [21], it was noted that the parity measurement-based encoding and decoding procedures can be implemented in constant depth regardless of architecture size. Since the decoding procedure corresponds to measuring vertices in one partition of a bipartite graph, it is clear that all measurements can be done simultaneously and corrected for in the other partition. The encoding procedure can be understood as measuring ancilla vertices of a larger graph state in the $X$-basis, which in particular produces the required bipartite graph (see e.g., [10, 11] for a characterisation of graph state deformations under Pauli measurements). It is known that all Pauli-measurements can be performed simultaneously in MBQC [18].
Secondly, there are a number of potential avenues for future research enabled by the results presented here. In the MBQC literature, there exist multi-particle entanglement purification protocols for bipartite graph states which exhibit favourable error thresholds for realistic scenarios [27, 28, 29]. Having demonstrated the prevalence of bipartite graph states in the parity framework, similar techniques may be of benefit for error mitigation in near-term implementations of parity quantum computations. Furthermore, proposals for fault-tolerant MBQC [30, 31, 32, 33] and universal blind quantum computation [34] could provide the foundation for fault-tolerant and cryptographic implementations of the parity framework. Conversely, developments related to quantum optimisation within the parity framework [35, 36, 37] could inspire similar developments in MBQC where application to optimisation problems remains relatively unexplored. A number of these avenues are already being pursued in separate work.
Acknowledgements.
We would like to thank B. Klaver and A. Messinger for useful discussions. This research was funded in whole or in part by the Austrian Science Fund (FWF) through the DK-ALM W1259125912591259-N27272727 and the SFB BeyondC F7102. For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. This work was also co-funded by the European Union (ERC, QuantAI, Project No. $101055129$). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.